OpenStudy (anonymous):

Given the equation y − 3 = 1/2(x + 6) in point-slope form, identify the equation of the same line in standard form. Can someone please help me?

OpenStudy (anonymous):

solve for \(y\) using the following steps 1) distribute the \(\frac{1}{2}\) on the right 2) add 3 to both sides

OpenStudy (anonymous):

I got x-2y= -12

OpenStudy (anonymous):

oh i should learn to read, hold on

OpenStudy (anonymous):

yes, you are right my instructions were wrong good work!

OpenStudy (anonymous):

Thank you! Can you help me with one more?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

not that i actually helped you with that one this time i will try to be more careful

OpenStudy (anonymous):

I try my hardest lol, sometimes I just get stumped, ok, here it goes.

OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the point (6, −3) and has a slope of 1/2.

OpenStudy (anonymous):

I am copying the answers

OpenStudy (anonymous):

y − 6 = one half(x + 3) y = one halfx − 6 y + 3 = one half(x − 6) x − 2y = 12

OpenStudy (anonymous):

k this one is a straight forward substitution in to the point - slope formula \[y-y_1=m(x-x_1)\] replace \(m\) by \(\frac{1}{2}\) replace \(x_1\) by \(6\) and replace \(y_1\) by \(-3\)

OpenStudy (anonymous):

I am so sorry, but I still do not understand. Some things in this lesson comes to me, and others, not so much.

OpenStudy (anonymous):

do you know what i mean by "make a direct substitution"?

OpenStudy (anonymous):

I will assume to make an automatic replacement?

OpenStudy (anonymous):

yes so where you see \(m\) put \(\frac{1}{2}\) for example

OpenStudy (anonymous):

\[\large y-\color{red}{y_1}=\color{green}m(x-\color{blue}{x_1})\] \[\color{green}m=\color{green}{\frac{1}{2}}\\ \color{blue}{x_1}=\color{blue}6\\ \color{red}{y_1}=\color{red}{-3}\]

OpenStudy (anonymous):

make the substitutions and let me know what you get

OpenStudy (anonymous):

My exam just reset so I lost everything..... Well, thank you.

OpenStudy (anonymous):

everything? damn!

OpenStudy (anonymous):

I know, not to mention all the questions changed.

OpenStudy (anonymous):

hurry!

OpenStudy (anonymous):

Ok for the question: Given the equation y − 4 = three fourths(x + 8) in point-slope form, identify the equation of the same line in standard form.

OpenStudy (anonymous):

I got 3x-4y= -40

OpenStudy (anonymous):

Would that be correct?

OpenStudy (anonymous):

that is what i get yes

OpenStudy (anonymous):

OK, hang on, I am going to try more lol

OpenStudy (anonymous):

k

OpenStudy (anonymous):

Ok so the question is: Choose the point-slope form of the equation below that represents the line that passes through the point (−1, 6) and has a slope of −3. y − 6 = −3x − 3 y − 6 = −3(x + 1) y = −3x + 3 3x + y = 3

OpenStudy (anonymous):

And I got Y-6= -3 (x+1)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Woo! Ok, I have more, sorry to bug you, I just don't want to mess this up.

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

Given the equation y − 4 = three fourths(x + 8) in point-slope form, identify the equation of the same line in slope-intercept form. −three fourthsx + y = 10 3x − 4y = −40 y = three fourthsx + 12 y = three fourthsx + 10

OpenStudy (anonymous):

And I got 3x-4y= -10

OpenStudy (anonymous):

that is standard form slope intercept form looks like \(y=mx+b\)

OpenStudy (anonymous):

Hmm, Ok lemme figure this out.

OpenStudy (anonymous):

ok let me know if you get stuck

OpenStudy (anonymous):

Y= 3/4x + 10?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

OK, just one more lol

OpenStudy (anonymous):

k no problem

OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the points (−6, 4) and (2, 0). y − 4 = −one half(x + 6) y − 4 = 2(x + 6) y + 6 = −one half(x − 4) y + 6 = 2(x − 4)

OpenStudy (anonymous):

And I got Y- 4 =2 (x+6)

OpenStudy (anonymous):

no your slope is wrong

OpenStudy (anonymous):

Poop, ok, let me try to fix this.

OpenStudy (anonymous):

\[\frac{0-4}{2+6}=\frac{-4}{8}=-\frac{1}{2}\]

OpenStudy (anonymous):

oops sorry i spoiled it

OpenStudy (anonymous):

Lol so it would be Y-4= -1/2 (x+6)?

OpenStudy (anonymous):

Sorry, I'm trying lol

OpenStudy (anonymous):

yes that is it

OpenStudy (anonymous):

you are doing fine

OpenStudy (anonymous):

Thank you! I got a 100% on my quiz :D

OpenStudy (anonymous):

yay

OpenStudy (anonymous):

you realize except for the last one you did all the work, i just checked it

OpenStudy (anonymous):

Yeah, I am proud of myself. But still, thank you for checking :)

OpenStudy (anonymous):

yw